OpenStudy (anonymous):

How do i change this equation x^2 + y^2 - 8y=20 Into this form (x-h)^2 + (y-k)^2= r^2 I am graphing a circle Please help!!!!

OpenStudy (dan815):

complete square for the y variable

OpenStudy (dan815):

x^2+y^2-8y+16=20+16 x^2+(y-4)^2=36

OpenStudy (dan815):

x^2+(y-4)^2=6^2

OpenStudy (dan815):

(x-0)^2+(y-4)^2=6^2

OpenStudy (anonymous):

x=\[\sqrt{20-y ^{2+8y, -\sqrt{20y ^{2+8y}}}}\]

OpenStudy (anonymous):

Thankyou @dan815